OpenStudy (anonymous):

Revenue is given by R(q)=444q and cost is given by C(q)=10000+3q^2 a) At what quantity is profit maximized? b) What is the profit at this production level?

OpenStudy (anonymous):

Ah, I figured it out. Profit = Revenue - Cost P(q)=R(q)-C(q) P(q)=444q - (10000 + 3q^2) P(q) = 3q^2 + 444q -10000 a) P'(q) = 555-6q 444= -6q |q|= 74 b) P(74) = -3(74)^2 + 444(74) - 10000 = $6425